Simplify the following:
5-3(x+2)
4(x-z)+5(y-x)

Answer by ewatrrr(24785)

5-3(x+2) = 5 - 3x - 6 = -(3x+1)
4(x-z)+5(y-x) = 4x - 4z + 5y - 5x = 5y - x - 4z